Given: angle 1 and angle 2 form a linear pair
angle 2 + angle 3 + angle 4 = 180
Prove: angle 1 = angle 3 and angle4

Answer by solver91311(24713)

Linear pairs of angles are supplementary, meaning that the sum of their measures is 180.

That means that a1 + a2 = 180, which we can write as a1 = 180 - a2.

But we also know that a2 + a3 + a4 = 180, which we can write as a3 + a4 = 180 - a2.

Since a = b and c = b if and only if a = c, a3 + a4 = a1.